The acceleration of a particle is increasing linearly with time t as bt. The particle starts from origin with initial velocity u. The distance travelled by the particle in time t will be:

 

a) ut+(1/3)bt2

 

b) ut+(1/2)bt2

 

c) ut+(1/6)bt3

 

d) ut+(1/3)bt3

 

Grade:11

1 Answers

Vikas TU
7 years ago
Displacement is the integral of accelration twice if the relation is gien in time.
Therefore,
a = bt (given)
dv/dt = bt
dv = bt dt
integerating both sides we get,
v = bt^2/2
dx = bt^2/2 dt
x = bt^3/6
Total distance = > ut +  (1/6)bt^3 m.